OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, -1) and has a slope of 5.

OpenStudy (anonymous):

y - 1 = 5(x - 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y - 2 = 5(x - 1)

OpenStudy (anonymous):

Equation with one point and slope is given by: \[\large y - y_1 = m(x - x_1)\] \[(x_1, y_1) = (-2,-1)\] \[m = 5\] Just plug in the values and find the equation..

OpenStudy (anonymous):

whats the answer>?

OpenStudy (anonymous):

Do you only want the answer ??? And not the solution ??

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

I asked something to you.. Please reply.. Do you only want the answer ??

OpenStudy (anonymous):

Answer is quite simple.. It is one of the four choices you have written above.. Good Luck..

OpenStudy (anonymous):

so what is the answer

jabez177 (jabez177):

@just_one_last_goodbye